Question 4 The value of sin 4 is 4 points оооо Question 5 The value of sin(-210°) is
Answer #1

sin ( ) - sin(ā + a) – - sin(a)​​​​​​

Sin (-210) = - sin(210) = -sin(180 +30) = -(-1/2)= 1/2